OpenStudy (anonymous):

Which of the following equations would graph a line parallel to 3x -y=7? A. y=7x-3 B. y=-3x+7 C. y=-7x+3 D. y=3x

OpenStudy (anonymous):

do you know what slope means first of all?

OpenStudy (anonymous):

Yes, I know the equation to find slope.

OpenStudy (anonymous):

place the equation in point-intercept form y=mx+b then find the equation that has the same m (which is parallel to the line)

OpenStudy (anonymous):

no not to find slope the meaning of slope

OpenStudy (anonymous):

then know the meaning of what parallel is

OpenStudy (anonymous):

I know what slope and parallel mean, but I can't put it into words.

OpenStudy (mathlegend):

y = mx+b Can you locate the slope in this formula?

OpenStudy (anonymous):

m = slope

OpenStudy (mathlegend):

Put this: 3x -y=7 in slope intercept form

OpenStudy (anonymous):

y=3x+7 Is this correct?

OpenStudy (mathlegend):

Okay when we moved everything to the right... -y= -3x +7 Now, when you divide by a negative don't all your signs change @babysmith ? So double check that...

OpenStudy (anonymous):

i mean you do not need to put slope and parallel into words just know what they mean in saying parallel means their slope must be the same

OpenStudy (anonymous):

Okay, then yes I understand that.

OpenStudy (mathlegend):

So once you divided by that -1... you should have gotten what @babysmith ?

OpenStudy (anonymous):



OpenStudy (mathlegend):

@babysmith -y= -3x +7 Can you solve this? Capturing your earlier mistake?

OpenStudy (anonymous):

I don't understand what I'm supposed to divide for. I thought I would leave the equation like this.

OpenStudy (mathlegend):

You should never have a -y... it should always be positive so divide by negative on both sides.. and tell me what you get.

OpenStudy (anonymous):

y = x+7 ?

OpenStudy (mathlegend):

\[\frac{ -y }{ -1 }=\frac{ -3x }{ -1 }-\frac{ 7 }{ -1 }\]

OpenStudy (mathlegend):

Try again

OpenStudy (anonymous):

y = 3x - 7 ?

OpenStudy (mathlegend):

It should be plus 7... like you had before. Just make sure you keep in mind every single sign. y = 3x + 7 They both had double negatives...

OpenStudy (mathlegend):

Being with what @iforgot told you.. that a parallel line would have the same slope... which answer would you choose?

OpenStudy (anonymous):

Ohhh okay I didn't notice the subtractions sign.

OpenStudy (mathlegend):

Any ideas @babysmith ?

OpenStudy (anonymous):

B.

OpenStudy (mathlegend):

The sign does matter @babysmith

hero (hero):

Actually, y = 3x - 7 is correct.

hero (hero):

Parallel lines have the same slope. So whatever the answer is, it must have the same slope as y = 3x - 7. @babysmith, what is the slope of y = 3x - 7 ?

OpenStudy (anonymous):

3.

OpenStudy (anonymous):

Oh my goodness that was so simple. I didn't even realize

hero (hero):

So which answer choice has the same slope?

OpenStudy (anonymous):

D!

hero (hero):

There we go. We all have to watch our signs. They are the cause of much trouble in mathematics if we mix them up. Even I myself have to watch it.

OpenStudy (anonymous):

Thank you all! (I never had a problem in mathematics until I started online school, fyi.)